# On the adjoint Selmer groups of semi-stable elliptic curves and Flach’s zeta elements

### v4

@@ 6. Proofs of corollaries
This follows from the tautological bound in §5.5 and the constant multiple $\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}=\dfrac{p^{\lambda}}{\mathrm{deg}(\phi)}$.
@@
The equivalence between (1) and (2) follows from the last argument in §$\S$5.5. The equivalence between the Bloch–Kato conjecture (2) and the modularity lifting theorem (3) follows from Wiles’ numerical criterion [DDT97, Thm. 5.3]. See also [DFG04].
@@
$\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell)\right)=\mathrm{ord}_{p}\left(c(\ell)\right)=0.$
@@ Acknowledgement
We thank Francesc Castella, Henri Darmon, Matthias Flach, Minhyong Kim, Masato Kurihara, Tom Weston, and Christopher Skinner for their interests in our work. We are benefited from the discussion with Ashay Burungale, Haruzo Hida, Antonio Lei, Gyujin Oh, and Ryotaro Sakamoto. We would like to thank Marco Sangiovanni Vincentelli for kindly sharing [SS] with us and for his interest on the cyclicity result.

### v6

@@ 6. Proofs of corollaries
This follows from the tautological bound in §5.5 and the constant multiple $\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}=\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\mathrm{deg}(\phi)}$.
@@
The equivalence between (1) and (2) follows from the last argument in $\S$5.5. The equivalence between the Bloch–Kato conjecture (2) and the modularity lifting theorem (3) follows from Wiles’ numerical criterion [DDT97, Thm. 5.3]. See also [DFG04].
@@
$\mathrm{ord}_{p}\left(\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell)\right)=\mathrm{ord}_{p}\left(c(\ell)\right)=0.$
@@ Acknowledgement
We thank Francesc Castella, Henri Darmon, Matthias Flach, Minhyong Kim, Masato Kurihara, Tom Weston, and Christopher Skinner for their interests in our work. We are benefited from the discussion with Ashay Burungale, Haruzo Hida, Antonio Lei, Gyujin Oh, and Ryotaro Sakamoto. We would like to thank Marco Sangiovanni Vincentelli for kindly sharing [SS] with us and for his interest on the cyclicity result. We sincerely thank the referees for the thorough reading and helpful comments, which have greatly improved the clarity of our exposition and have removed errors in an earlier version. In particular, the computation of the constant multiple is significantly simplified by the suggestion of a referee.